///Apiary Record : 10th July 2016.

Apiary Record : 10th July 2016.

Still under EFB shutdown until being rechecked by the SBI on the 19th July.  So quick check through, making sure no colonies were about to swarm and that they had enough room,stores and that they had no sign of any disease.
Weather:   Sunny after earlier rain/drizzle.  19 degrees C,  Breezy.
Hive#1   Inspected by Liz,    Brood Chamber  14 x 12,  QNS,  No Qc’s,  No Brood, No eggs, stores/room ok.   1 super  75% full,   Queenless..  Introduced a frame with 2 QC’s from 15H( soon to be the new Colony 3).  Drone Trap removed,  far too many drones in the hive. leave 4 weeks.  Temper : 9
Hive#2   Not Checked as sealed QC left on f2 on 3/7/16    will be checked on 19/7 by SBI
Hive#4  Inspected by Liz.   Brood chamber -National Deep.  QNS,  2 Qc’s Brood  AS,  and new day old eggs,  2 Supers  nearly full, stores/room ok.   –  Knocked down both charged Qc’s.  Temper: 7
Hive#5  Not gone through as poss virgin Queen.
Hive#6  Inspected by Liz,    Brood chamber,  Deep and a Shallow.   QNS, possibly swarmed.        4 charged Qc ‘s   Brood  5D, 4S,  2 supers, top full, –  knocked down 2 Qc’s leaving 1 sealed, and 1 charged but open, on nframe with red pin.  No eggs seen at all.-  leave for 4 weeks.                   Temper:  6.5
Hive#7  Inspected by Liz,   Brood chamber: Deep.   QNS,  no Qc’s, stores/room ok,  2 supers.     new eggs in drone comb only, so may be a drone laying queen, Combined with Nuc 16A as 2 Sealed Qc’s on f5.
Hive#7A   Inspected by Richard,    Brood chamber:  4 Shallows.   QNS, no Qc’s.  7 frames of brood, all in top super. some 6 -7 day old larva,  no new eggs.  Q gone?  Dirty comb in bottom supers needs replacing.
Hive#8    Inspected by Richard,    Brood chamber : Deep.   QNS, 2 charged Qc’s,  6 f BAS, new eggs, 2 supers, top full.  stores/room ok,   Charged Qc’s taken down.  lots of bees looks like they may swarm.
Hive#15F  Inspected by Liz.     Brood chamber: Deep.  QS,  no Qc’s.  6 f BAS,  stores/room ok,     1 super,  –  Knocked down several play cups.
Hive#15H   Inspected by Richard.   Brood chamber: Deep.   QS,  2 QC’s ,  7 f BAS and new eggs.  Stores/room ok,  1 super . Part filled,-   Frame with 2 Qc’s given to Hive 1.    Temper:  9
Nuc # 16A   Combined with Hive#7.   but no Q found, 2 sealed QC’s  left.
Nuc# 16B   Not gone through – though bees bringing in pollen
Nuc# 16C  Not gone through –  bees flying
 
 

By |2016-07-14T22:11:38+01:00July 14th, 2016|Apiary, 2016|Comments Off on Apiary Record : 10th July 2016.